OpenStudy (anonymous):

Find f f ''(x)= -2+18x-12(x^2) f(0)=3 f '(0)=14 No idea how to do this. Please help!

OpenStudy (anonymous):

f'(x) = -2x + 9x^2 - 4(x^3) + A f(x) = -x^2 + 3x^3 - x^4 + Ax + B substitute values to find A and B

OpenStudy (anonymous):

oh so f(x)= -x^2 + 3x^3 -x^4 +14x + 3?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that is correct.

OpenStudy (anonymous):

Thanks!!